# Eccentric Mergers in AGN Discs: Influence of the Supermassive Black-Hole on Three-body Interactions

### v1

@@ 4.4 Topology of end-states
We only show a portion of the overall analyzed parameter space to look more closely at what happens when the SMA approaches the size of $R_{\rm H}$. From the phase space distribution, a few elements stand out. In terms of $\phi_{0}$, for both the CR and CT cases the phase-space structure has a trivial periodicity, where the pattern repeats but in opposite colors after $180^{\circ}$. In terms of the structure along the SMA, we see that as we approach the $R_{\rm H}$ limit, the regions widen and for some angular intervals, e.g. in the CR case (left), there is really only one outcome possible. For example, at $200^{\circ}$ in the CR case, the system can only result in outcome ‘BS13’ as the binary has no time to undergo significant orbital evolution before it encounters the single, which in turn turns into this restricted outcome space. At lower SMA, the binary can for just small changes in SMA undergo several revolutions, before the single arrives and initiates the interaction (see Eq. 31), which give rise to the seen band-like structure in both the CR and the CT case. Another interesting behavior that we further analyse in Fig. 8, is that in the CR case (left) there appears to be a significant higher fraction of non-chaotic regions than chaotic ones compared to the CT case (right). Intuitively, this is to be expected from the setup, as the single, which arrives from the Hill sphere, generally encounters the binary from ‘below’ (since the binary in this setup is close to the SMBH than the single), as seen in Fig. 3, and therefore initially orbits the binary in the same direction as the CT setup. The objects therefore encounter each other with a lower relative velocity, which generally leads to a temporary dynamical capture of the single. This effect starts the resonating interaction, which corresponds to the seed to enter a chaotic state. This is in contrast to the CR case, where the single encounters one of the binary objects with a relatively high velocity, which is more likely to result in a prompt exchange, which might bring the objects out of the Hill sphere after the first encounter. What is not expected is that the resonating interaction and corresponding chaotic regions are so rare in the CR case.
Furthermore, the phase space figures provide important hints as to how and where the eccentric BBH mergers form, which is also illustrated in Fig. 9. Here we highlight mergers formed in the chaotic regions (left side) and in the prompt regions (right side). The interactions taking place in the resonating and chaotic regions are much more prone to additional effects such as gas, if present, and are therefore likely be greatly affected, as they have much more time to lose energy through drag over the duration of the resonating state. The opposite can be stated regarding prompt mergers, although they could turn into resonating interactions if enough drag is acting on them before they leave the Hill sphere.
@@ 4.5 Out-of-plane interactions
Until now we have focused on properties and outcomes of the co-planar setup, where all four objects (SMBH, binary+single) are interacting in the same plane, as this configuration gives rise to the most distinct outcomes relative to the well studied isotropic case (Samsing et al., 2014; Samsing, 2018). However, the orbital evolution of objects inside the disc is believed to be subject to e.g. turbulence in the disc and mechanisms due to dynamical heating (Whitehead et al., 2023), which ultimately result in mutual interactions that are not co-planar, which we refer to as out-of-plane interactions. We quantify the out-of-plane interaction by the inclination angle, $i$, which refers to the angle between the incoming single relative to the orbital BBH and disc plane. As shown in the top part of Fig. 11, we place the single incoming BH at the Hill sphere at angle $i$ above the disc-plane, with a velocity vector that is equivalent to the co-planar setup, i.e. we consider the situation where the single BH initially has zero velocity in the direction perpendicular to the disc. A similar setup was explored in S22, but in their case the single BH was assumed to be coming from infinity in a plane with inclination $i$ relative to the BBH orbital plane, in which case the single BH passes through the Hill sphere approximately with the same angle $i$, but with a slightly different angular momentum. In S22 it was shown that the outcome probability of eccentric mergers strongly depends on $i$, with nearly one order-of-magnitude decrease when the angle has opened to $1^{\circ}$ relative to the $0^{\circ}$ co-planar case. This strong dependence on the orbital plane inclination angle naturally questions how the results actually depend on the assumption that the single BH approached from infinity as normally assumed in scattering systems. It could be that if the object is initiated at the Hill sphere, and not propagated to the Hill sphere from far away, the dependence on $i$ is much weaker, as what really matters is how much angular momentum the single BH brings. With this motivation, we now consider the results shown in Fig. 11, which shows the probability for eccentric mergers ($e>0.1$ with a peak frequency $f_{\rm gw}\geq 10$ Hz), resulting from our out-of-plane Setup I where the single starts at the Hill sphere at an angle $i$ above the plane. In Fig. 11 the BH masses are equal and set to $15\,M_{\odot}$. We show in total four different setups characterized by two different SMAs (1.7 AU, cross symbols and 1 AU, circles) and for CR (purple) and CT (orange) binaries. As seen, the general trend is consistent with S22, i.e. at angle $10^{-2}$ the probabilities converge to the value found in the co-planar case shown in Fig. 7, with a rapid decrease by almost 2-3 orders of magnitude once $i$ passes $1^{\circ}-10^{\circ}$. Even in this setup, the probability for eccentric mergers therefore still seems to be very sensitive to the inclination angle, which add further motivation for studying how the gas has an impact on the alignment of the objects in the disc. This is particularly important when considering the binaries surviving the interaction, as they are scattered both out of the disc-plane and have their orbital momentum tilted (S22). Many of these binaries are likely to undergo subsequent interactions inside the disc, but how aligned they are in the disc at the time of interaction depends on the alignment timescales relative to the interaction timescale, as well as the properties of the binary after interaction. The subsequent question is naturally what the expected out-of-plane angle could be in AGN-disc like environments. Assuming the binary orbital angular moment is perpendicular to the disc angular momentum, a simple estimate can be made in the case the Hill sphere is larger than the disc height, in which case the maximum opening angle is around the AGN-disc height divided by the Hill radius, i.e. $max(i)\approx h/R_{H}\approx(h/R)\times\left(M/m\right)^{1/3}$. The maximum opening angle can span a range of different values, depending on both aspect-ratio and mass ratio. For example, for a system described by $m=10^{1}M_{\odot}$, $M=10^{7}M_{\odot}$, then $max(i)\approx h/R_{\rm H}\times 10^{2}$, i.e. for thin-disc models with $h/R\sim 10^{-3}$ like Shakura & Sunyaev (1973), the characteristic angle could very well be $<1^{\circ}$. Reading from the bottom panel in Fig. 11, this would still mean that eccentric mergers are likely to form in over-abundance in disc environments compared to isotropic environments. However, for more realistic disc models, this simple estimate could vary throughout the disc, reaching values of $h/R\sim 1$ (e.g. Sirko & Goodman, 2003; Thompson et al., 2005). This structure would make some regions more prone to forming eccentric mergers than others.
@@ 5 Conclusions and caveats
  • Theory of 3-body interactions in a tidal field. When binary-single objects interact, they often undergo a large number of resonant states that are characterized by long excursions of the single relative to the binary. Each of these IMS states, have been shown to be one of the main pathways to form eccentric mergers in systems where such interactions can take place. In Sec. 2.2, we have extended the theory of how such eccentric 3-body mergers form in resonant interactions, by including the effect from a tidal field characterized by a Hill sphere boundary. By the use of statistical theory for how the energy and angular momentum is distributed in chaotic encounters, based on works by Valtonen & Karttunen (2006); Stone & Leigh (2019), we showed that the probability for 3-body mergers, i.e mergers forming during the interactions, generally are reduced by a factor of $(1-2a_{0}/R_{\rm H})^{2}$, where $a_{0}$ is the initial BBH semi-major axis and $R_{\rm H}$ is the Hill sphere (see Eq. 14). While true chaotic interactions are difficult to achieve in environments with a strong nearby perturber, we identify this as the first indication of how a nearby SMBH might impact the fraction of eccentric mergers. Fig. 5, shows this correction factor applied on the results from S22.
@@
  • PN-Simulations of eccentric mergers under the influence of the SMBH. We performed a systematic study of PN scatterings between a BBH and a single incoming BH taking place near the SMBH. In this setup, we especially explored the formation probability of eccentric BBH mergers ($e$ > 0.1 at $f_{\rm gw}$ > 10 Hz), and how this depends on the SMA of the initial BBH $(a_{0})$ relative to the effective Hill sphere of the SMBH ($R_{\rm H}$), its orbital angular momentum relative to its direction around the SMBH (co- and counter-rotating), the location of the single relative to the BBH (Setup I and Setup II, see Fig. 3), as well as the initial inclination between the BBH and single BH. In all scatterings we initiated the single BH at the Hill sphere, to emulate the likely process that takes place when a BBH and a single black hole meet each other in an AGN-disc environment through disc migration. S22 showed that three-body scatterings in disc-like configurations with no SMBH included, produce a high fraction of eccentric mergers. With more realistic setups we here find that eccentric mergers are indeed also found to be produced in large numbers even when the interactions take place close to the SMBH. The reason is partly that many of the near co-planar interactions produce mergers that are formed through prompt interactions (see Fig. 9), that are not greatly influenced by the SMBH upon formation. Finally, as shown in Fig. 6, many of the 3-body mergers do actually start with a GW peak frequency that can be far above 10 Hz, hinting that a significant fraction might show up as bursts or prompt mergers, rather than classical eccentric inspirals, where the system evolves from eccentric to circular over hundreds of cycles. This result is in agreement with the analysis of Trani et al. (2023) and Rom et al. (2023), where for the latter they explore the effects of tidal forces coming from the SMBH on binary capture in AGNs. These findings further motivate searches for burst-like GW signals.
@@
  • Importance of orbital inclinations. No interactions are expected to be perfectly co-planar, studying how our results depend on the initial inclination angle between the BBH and single when the single enters the Hill sphere (angle $i$ as seen in Fig. 11), is therefore of major importance. In Sec. 4.5, we explored the fraction of eccentric BBH mergers as a function of the initial inclination angle $i$, and found that across the different constrained setups, the probability do falls off quickly with $i$ as was also pointed out in S22. All AGN-disc environments are different from each other, but we argue that for thin disc models, in regions where the maximum opening angle is $i\sim 1^{\circ}$, a relatively high fraction of eccentric mergers are likely to be produced.
  • Potential impact of gas in the accretion disc. As discussed, we do not include gaseous effects in our analysis but acknowledge their potential impact on the presented results and on the overall fraction of eccentric mergers. A relevant question to be explored is if the additional gas drag force term keeps the objects contained in the Hill sphere, ultimately leading to more chaotic interactions in the co-rotating and counter-rotating case as a contrast to the high fraction of prompt interaction regions we observe from e.g. Fig. 8 & 9. In addition, a full dynamical model of migrating BHs inside AGN-disc-like environments is necessary to state what relevant values of $R_{\rm H}$ and SMA are to be expected, and therefore how many interactions that can be considered chaotic or not. As far as the nature and outlook of mergers is concerned, the scattering seen in the bottom panel of Fig. 4 might in comparison not be much affected by gas, as the binary undergoing merger is formed promptly. Therefore, we speculate that gas might not have a great impact on the formation of such prompt mergers, which could hint that our results might also hold when gaseous effects are included. There are however, several reasons to believe that our considerations and assumptions from above need to be adjusted, as the relevant setup that we consider involves binaries that encounter singles as a result of migration through an AGN-disc (e.g. Secunda et al., 2019, 2020; Tagawa et al., 2020). Lastly, weather gas friction is expected to damp or pump the eccentricity of the binary is a non-trivial answer, as its evolution depends on a number of factors of the parameter space, such as mass ratio or whether binary is in prograde or retrograde motion with respect to the gas (e.g. Li & Lai, 2022; Li & Lai, 2023b; Li & Lai, 2023a; Dittmann et al., 2023; Valli et al., 2024).
@@ Acknowledgements
GF thanks the GW-Astro group at NBIA, in particular Martin Pessah and Daniel J. D’Orazio, for insightful discussions and constant support, as well as the TIDY-NYC group for interesting discussions. The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for performing the simulations presented in the paper. This work was funded by the Villum Fonden Grant No. 29466 and ERC Starting Grant no. 101043143.

### v2

@@ 4.4 Topology of end-states
We only show a portion of the overall analyzed parameter space to look more closely at what happens when the SMA approaches the size of $R_{\rm H}$. From the phase space distribution, a few elements stand out. In terms of $\phi_{0}$, for both the CR and CT cases the phase-space structure has a trivial periodicity, where the pattern repeats but in opposite colors after $180^{\circ}$. In terms of the structure along the SMA, we see that as we approach the $R_{\rm H}$ limit, the regions widen and for some angular intervals, e.g. in the CR case (left), there is really only one possible outcome. For example, at $200^{\circ}$ in the CR case, the system can only result in outcome ‘BS13’ as the binary has no time to undergo significant orbital evolution before it encounters the single, which in turn turns into this restricted outcome space. At lower SMA, the binary can for just small changes in SMA undergo several revolutions, before the single arrives and initiates the interaction (see Eq. 31), which give rise to the seen band-like structure in both the CR and the CT case. Another interesting behavior that we further analyse in Fig. 8, is that in the CR case (left) there appears to be a significant higher fraction of non-chaotic regions than chaotic ones compared to the CT case (right). Intuitively, this is to be expected from the setup, as the single, which arrives from the Hill sphere, generally encounters the binary from ‘below’ (since the binary in this setup is close to the SMBH than the single), as seen in Fig. 3, and therefore initially orbits the binary in the same direction as the CT setup. The objects therefore encounter each other with a lower relative velocity, which generally leads to a temporary dynamical capture of the single. This effect starts the resonating interaction, which corresponds to the seed to enter a chaotic state. This is in contrast to the CR case, where the single encounters one of the binary objects with a relatively high velocity, which is more likely to result in a prompt exchange, which might bring the objects out of the Hill sphere after the first encounter. What is not expected is that the resonating interaction and corresponding chaotic regions are so rare in the CR case.
Furthermore, the phase space figures provide important hints as to how and where the eccentric BBH mergers form, which is also illustrated in Fig. 9. Here we highlight mergers formed in the chaotic (left side) and in the prompt regions (right side). The interactions taking place in the resonating and chaotic regions are much more prone to additional effects such as gas, if present, and are therefore likely to be greatly affected, as they have much more time to lose energy through drag over the duration of the resonating state. The opposite can be stated regarding prompt mergers, although they could turn into resonating interactions if enough drag is acting on them before they leave the Hill sphere.
@@ 4.5 Out-of-plane interactions
Until now we have focused on properties and outcomes of the co-planar setup, where all four objects (SMBH, binary+single) are interacting in the same plane, as this configuration gives rise to the most distinct outcomes relative to the well studied isotropic case (Samsing et al., 2014; Samsing, 2018). However, there are different factors that can result in a non-fully 2-D distribution of BHs embedded in the disc. When BHs in the nuclear star cluster are captured by the disc due to drag force processes they do not necessarily get aligned in fully co-planar orbits (e.g. Fabj et al., 2020; Nasim et al., 2023; Wang et al., 2024), dynamical heating will occur if a larger population of the disc is present as discussed in Stone et al. (2017), and the orbital evolution of objects inside the disc is believed to be subject to e.g. gas turbulence (Whitehead et al., 2024). All these processes ultimately result in mutual interactions that are not co-planar, which we refer to as out-of-plane interactions.
@@
We quantify the out-of-plane interaction by the inclination angle, $i$, which refers to the angle between the incoming single relative to the orbital BBH and disc plane. As shown in the top part of Fig. 11, we place the single incoming BH at the Hill sphere at angle $i$ above the disc-plane, with a velocity vector that is equivalent to the co-planar setup, i.e. we consider the situation where the single BH initially has zero velocity in the direction perpendicular to the disc. A similar setup was explored in S22, but in their case the single BH was assumed to be coming from infinity in a plane with inclination $i$ relative to the BBH orbital plane, in which case the single BH passes through the Hill sphere approximately with the same angle $i$, but with a slightly different angular momentum. In S22 it was shown that the outcome probability of eccentric mergers strongly depends on $i$, with nearly one order-of-magnitude decrease when the angle has opened to $1^{\circ}$ relative to the $0^{\circ}$ co-planar case. This strong dependence on the orbital plane inclination angle naturally questions how the results actually depend on the assumption that the single BH approached from infinity as normally assumed in scattering systems. It could be that if the object is initiated at the Hill sphere, and not propagated to the Hill sphere from far away, the dependence on $i$ is much weaker, as what really matters is how much angular momentum the single BH brings. With this motivation, we now consider the results shown in Fig. 11, which shows the probability for eccentric mergers ($e>0.1$ with a peak frequency $f_{\rm gw}\geq 10$ Hz), resulting from our out-of-plane Setup I where the single starts at the Hill sphere at an angle $i$ above the plane. In Fig. 11 the BH masses are equal and set to $15\,M_{\odot}$. We show in total four different setups characterized by two different SMAs (1.7 AU, cross symbols and 1 AU, circles) and for CR (purple) and CT (orange) binaries. As seen, the general trend is consistent with S22, i.e. at angle $10^{-2}$ the probabilities converge to the value found in the co-planar case shown in Fig. 7, with a rapid decrease by almost 2-3 orders of magnitude once $i$ passes $1^{\circ}-10^{\circ}$. Even in this setup, the probability for eccentric mergers therefore still seems to be very sensitive to the inclination angle, which adds further motivation for studying how the gas has an impact on the alignment of the objects in the disc. This is particularly important when considering the binaries surviving the interaction, as they are scattered both out of the disc-plane and have their orbital momentum tilted (S22). Many of these binaries are likely to undergo subsequent interactions inside the disc, but how aligned they are in the disc at the time of interaction depends on the alignment timescales relative to the interaction timescale, as well as the properties of the binary after interaction. The subsequent question is naturally what the expected out-of-plane angle could be in AGN-disc like environments. Assuming the binary orbital angular moment is perpendicular to the disc angular momentum, a simple estimate can be made in the case the Hill sphere is larger than the disc height, in which case the maximum opening angle is around the AGN-disc height divided by the Hill radius, i.e. $max(i)\approx h/R_{H}\approx(h/R)\times\left(M/m\right)^{1/3}$. The maximum opening angle can span a range of different values, depending on both aspect-ratio and mass ratio. For example, for a system described by $m=10^{1}M_{\odot}$, $M=10^{7}M_{\odot}$, then $max(i)\approx h/R_{\rm H}\times 10^{2}$, i.e. for thin-disc models with $h/R\sim 10^{-3}$ like Shakura & Sunyaev (1973), the characteristic angle could very well be $<1^{\circ}$. Reading from the bottom panel in Fig. 11, this would still mean that eccentric mergers are likely to form in over-abundance in disc environments compared to isotropic environments. However, for more realistic disc models, this simple estimate could vary throughout the disc, reaching values of $h/R\sim 0.1$ (e.g. Sirko & Goodman, 2003; Thompson et al., 2005). This structure would make some regions more prone to forming eccentric mergers than others.
@@ 5 Conclusions and caveats
  • Theory of 3-body interactions in a tidal field. When binary-single objects interact, they often undergo a large number of resonant states that are characterized by long excursions of the single relative to the binary. Each of these IMS states has been shown to be one of the main pathways to form eccentric mergers in systems where such interactions can take place. In Sec. 2.2, we have extended the theory of how such eccentric 3-body mergers form in resonant interactions, by including the effect from a tidal field characterized by a Hill sphere boundary. By the use of statistical theory for how the energy and angular momentum is distributed in chaotic encounters, based on works by Valtonen & Karttunen (2006); Stone & Leigh (2019), we showed that the probability for 3-body mergers, i.e. mergers forming during the interactions, are generally reduced by a factor of $(1-2a_{0}/R_{\rm H})^{2}$, where $a_{0}$ is the initial BBH semi-major axis and $R_{\rm H}$ is the Hill sphere (see Eq. 14). While true chaotic interactions are difficult to achieve in environments with a strong nearby perturber, we identify this as the first indication of how a nearby SMBH might impact the fraction of eccentric mergers. Fig. 5, shows this correction factor applied on the results from S22.
@@
  • PN-Simulations of eccentric mergers under the influence of the SMBH. We performed a systematic study of PN scatterings between a BBH and a single incoming BH taking place near the SMBH. In this setup, we especially explored the formation probability of eccentric BBH mergers ($e$ > 0.1 at $f_{\rm gw}$ > 10 Hz), and how this depends on the SMA of the initial BBH $(a_{0})$ relative to the effective Hill sphere of the SMBH ($R_{\rm H}$), its orbital angular momentum relative to its direction around the SMBH (co- and counter-rotating), the location of the single relative to the BBH (Setup I and Setup II, see Fig. 3), as well as the initial inclination between the BBH and single BH. In all scatterings we initiated the single BH at the Hill sphere, to emulate the likely process that takes place when a BBH and a single black hole meet in an AGN-disc environment through disc migration. S22 showed that three-body scatterings in disc-like configurations with no SMBH included, produce a high fraction of eccentric mergers. With more realistic setups we here find that eccentric mergers are indeed also found to be produced in large numbers even when the interactions take place close to the SMBH. The reason is partly that many of the near co-planar interactions produce mergers that are formed through prompt interactions (see Fig. 9), that are not greatly influenced by the SMBH upon formation. Finally, as shown in Fig. 6, many of the 3-body mergers do actually start with a GW peak frequency that can be far above 10 Hz, hinting that a significant fraction might show up as bursts or prompt mergers, rather than classical eccentric inspirals, where the system evolves from eccentric to circular over hundreds of cycles. This result is in agreement with the analysis of Trani et al. (2023) and Rom et al. (2023), where for the latter they explore the effects of tidal forces coming from the SMBH on binary capture in AGNs. These findings further motivate searches for burst-like GW signals.
@@
  • Importance of orbital inclinations. No interactions are expected to be perfectly co-planar, studying how our results depend on the initial inclination angle between the BBH and single when the single enters the Hill sphere (angle $i$ as seen in Fig. 11), is therefore of major importance. In Sec. 4.5, we explored the fraction of eccentric BBH mergers as a function of the initial inclination angle $i$, and found that across the different constrained setups, the probability does falls off quickly with $i$ as was also pointed out in S22. All AGN-disc environments are different from each other, but we argue that for thin disc models, in regions where the maximum opening angle is $i\sim 1^{\circ}$, a relatively high fraction of eccentric mergers are likely to be produced.
  • Potential impact of gas in the accretion disc. As discussed, we do not include gaseous effects in our analysis but acknowledge their potential impact on the presented results and on the overall fraction of eccentric mergers. A relevant question to be explored is if the additional gas drag force term keeps the objects contained in the Hill sphere, ultimately leading to more chaotic interactions in the co-rotating and counter-rotating case as a contrast to the high fraction of prompt interaction regions we observe from e.g. Fig. 8 & 9. In addition, a full dynamical model of migrating BHs inside AGN-disc-like environments is necessary to state what relevant values of $R_{\rm H}$ and SMA are to be expected, and therefore how many interactions can be considered chaotic or not. As far as the nature and outlook of mergers is concerned, the scattering seen in the bottom panel of Fig. 4 might in comparison not be much affected by gas, as the binary undergoing merger is formed promptly. Therefore, we speculate that gas might not have a great impact on the formation of such prompt mergers, which could hint that our results might also hold when gaseous effects are included. There are however, several reasons to believe that our considerations and assumptions from above need to be adjusted, as the relevant setup that we consider involves binaries that encounter singles as a result of migration through an AGN-disc (e.g. Secunda et al., 2019, 2020; Tagawa et al., 2020). Lastly, whether gas friction is expected to damp or pump the eccentricity of the binary is a non-trivial answer, as its evolution depends on a number of factors of the parameter space, such as mass ratio or whether binary is in prograde or retrograde motion with respect to the gas (e.g. Li & Lai, 2022; Li & Lai, 2023b; Li & Lai, 2023a; Dittmann et al., 2023a; Valli et al., 2024).
@@ Acknowledgements
The authors thank the anonymous referee for the useful suggestions and comments. GF thanks the GW-Astro group at NBIA, in particular Martin Pessah and Daniel J. D’Orazio, for insightful discussions and constant support, as well as the TIDY-NYC group for interesting discussions. The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for performing the simulations presented in the paper. This work was funded by the Villum Fonden Grant No. 29466 and ERC Starting Grant no. 101043143.